### v2

[…]
1 Introduction
Markov decision process (MDP) (Puterman, 2014) is one of the most popular frameworks to model sequential decision-making by a single agent, with recent applications to game solving and reinforcement learning (Sutton and Barto, 2018; Mnih et al., 2013), healthcare decision-making (Bennett and Hauser, 2013; Steimle and Denton, 2017) and finance (Bäuerle and Rieder, 2011). Robust Markov decision processes (RMDPs) is a generalization of MDPs, dating as far back as Satia and Lave (1973) and extensively studied more recently after the seminal papers of Iyengar (2005) and Nilim and Ghaoui (2005). Robust MDPs consider a single agent repeatedly interacting with an environment with unknown instantaneous rewards and/or transition probabilities. The unknown parameters are assumed to be chosen adversarially from an uncertainty set, modeling the set of all plausible values of the parameters.
Most of the robust MDP literature has focused on computational and algorithmic considerations for the discounted return, where the future instantaneous rewards are discounted with a discount factor $\gamma\in(0,1)$. In this case, RMDPs become tractable under certain rectangularity assumptions, with models such as sa-rectangularity (Iyengar, 2005; Nilim and Ghaoui, 2005), s-rectangularity (Le Tallec, 2007; Wiesemann et al., 2013), k-rectangularity (Mannor et al., 2016), and r-rectangularity (Goh et al., 2018; Goyal and Grand-Clément, 2022). For rectangular RMDPs, iterative algorithms can be efficiently implemented for various distance- and divergence-based uncertainty sets (Givan et al., 1997; Iyengar, 2005; Ho et al., 2021; Grand-Clement and Kroer, 2021; Behzadian et al., 2021), as well as gradient-based algorithms (Wang et al., 2022; Kumar et al., 2023; Li et al., 2022, 2023).
Robust MDPs have found some applications in healthcare (Zhang et al., 2017; Goh et al., 2018; Grand-Clément et al., 2023), where the set of states represents the potential health conditions of the patients, actions represent the medical interventions, and it is critical to account for the potential errors in the estimated transition probabilities representing the evolution of the patient’s health, and in inverse reinforcement learning (Viano et al., 2021; Chae et al., 2022) for imitations that are robust to shifts between the learner’s and the experts’ dynamics. In some applications, introducing a discount factor can be seen as a modeling choice, e.g. in finance (Deng et al., 2016). In other applications like game-solving (Mnih et al., 2013; Brockman et al., 2016), the discount factor is merely introduced for algorithmic purposes, typically to ensure the convergence of iterative algorithms or to reduce variance (Baxter and Bartlett, 2001), and it may not have any natural interpretation. Additionally, large discount factors may slow the convergence rate of the algorithms. Average optimality, i.e., optimizing the limit of the average of the instantaneous rewards obtained over an infinite horizon, and Blackwell optimality, i.e., finding policies that remain discount optimal for all discount factors sufficiently close to $1$, provide useful objective criteria for environments with no natural discount factor, unknown discount factor, or large discount factor. To the best of our knowledge, the literature on robust MDPs that address these optimality notions is scarce. The authors in Tewari and Bartlett (2007); Wang et al. (2023) study RMDPs with the average return criterion but only focus on computing the optimal worst-case average returns among stationary policies, without any guarantee that average optimal policies may be chosen in this class of policies. Similarly, some previous papers show the existence of Blackwell optimal policies for sa-rectangular RMDPs, but they require some restrictive assumptions, such as polyhedral uncertainty sets (Tewari and Bartlett, 2007; Goyal and Grand-Clément, 2022) or unichain RMDPs with a unique average optimal policy (Wang et al., 2023).
Our goal in this paper is to study RMDPs with average and Blackwell optimality in the general case. Our main contributions can be summarized as follows.
  1. 1.
    Average optimality. For sa-rectangular robust MDPs with compact convex uncertainty sets, we show the existence of a stationary deterministic policy that is average optimal, thus closing an important gap in the literature. Additionally, we describe several strong duality results and we highlight that the worst-case transition probabilities need not exist, a fact that has been overlooked by previous work. We also discuss the case of robustness against history-dependent worst-case transition probabilities. In addition, we show for s-rectangular RMDPs that surprisingly, stationary policies may be suboptimal for the average return criterion. That is, history-dependent (Markovian) policies may be necessary to achieve the optimal average return.
  2. 2.
    Blackwell optimality. We provide an extensive treatment of Blackwell optimality for sa-rectangular RMDPs. In this case, we provide a counterexample where Blackwell optimal policies fail to exist, and we identify the pathological oscillating behaviors of the robust value functions as the key factor for this non-existence. We show however that approximate Blackwell optimal policies always exist. Finally, we introduce the notion of definable uncertainty sets, which are sets built on simple enough functions to ensure that the discounted value functions are “well-behaved”. The definability assumption captures virtually all uncertainty sets used in practice for sa-rectangular RMDPs, and we show that for sa-rectangular RMDPs with definable uncertainty sets, there always exists a stationary deterministic policy that is Blackwell optimal. We also highlight the connections between Blackwell optimal policies and average optimal policies. For s-rectangular RMDPs, we provide a simple example where Blackwell optimal policies do not exist, even in the simple case of polyhedral uncertainty.
  3. 3.
    Algorithms and numerical experiments. We introduce three algorithms that converge to the optimal average return, in the case of definable sa-rectangular uncertainty sets. We note that efficiently computing an average optimal policy for sa-rectangular RMDPs is a long-standing open problem in algorithmic game theory, and we present numerical experiments on three RMDP instances.
Overall, we provide a complete picture of average optimality and Blackwell optimality in sa-rectangular RMDPs, and we show surprising new results for s-rectangular RMDPs. Table 1 summarizes our main results and compares them with prior results for robust MDPs and nominal MDPs. Our results are highligted in bold.
Table 1: Properties of optimal policies for different objective criteria. The set of states and the set of actions are finite. Our results are in bold.
Uncertainty set $\mathcal{U}$
Discount
optimality
Average
optimality
Blackwell optimality
Singleton (MDPs)
stationary,
deterministic
stationary,
deterministic
stationary, deterministic
sa-rectangular,
compact, convex
stationary,
deterministic
stationary,
deterministic
  • may not exist
  • $\exists\;\pi$ stationary deterministic,
    $\pi\;\epsilon$-Blackwell optimal, $\forall\epsilon>0$
  • $\pi$ also average optimal
sa-rectangular,
compact, convex, definable
stationary,
deterministic
stationary,
deterministic
  • stationary, deterministic
  • $\pi$ also average optimal
s-rectangular,
compact convex
stationary,
randomized
history-dependent,
randomized
may not exist
Outline of the paper.
The rest of the paper is organized as follows. We introduce robust MDPs in Section 2 and provide our literature review there. We study average optimality in Section 3 and Blackwell optimality in Section 4. We study algorithms for computing average optimal policies for sa-rectangular RMDPs in Section 5 and we provide some numerical experiments in Section 6.
1.1 Notations.
Given a finite set $\Omega$, we write $\Delta(\Omega)$ for the simplex of probability distributions over $\Omega$, and $P(\Omega)$ for the power set of $\Omega$. We allow ourselves to conflate vectors in $\mathbb{R}^{|\Omega|}$ and functions on $\mathbb{R}^{\Omega}$. For any two vectors $\bm{u},\bm{v}\in\mathbb{R}^{\Omega}$, the inequality $\bm{u}\leq\bm{v}$ is understood componentwise: $u_{s}\leq v_{s},\forall\;s\in\Omega$, and for $\epsilon\geq 0$, we overload the notation $\bm{u}\leq\bm{v}+\epsilon$ to signify $u_{s}\leq v_{s}+\epsilon,\forall\;s\in\Omega$. We write $\bm{e}=(1,...,1)$ and its dimension depends on the context.
2 Robust Markov decision processes
In this section, we introduce discounted robust MDPs, and discuss the connections with stochastic games which prove useful in the rest of the paper.
2.1 Discount optimality
A robust Markov decision process is defined by a tuple $\left(\mathcal{S},\mathcal{A},\bm{r},\mathcal{U},\bm{p}_{0}\right)$ where $\mathcal{S}$ is the finite set of states, $\mathcal{A}$ is the finite set of actions available to the decision-maker, and $\bm{r}\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}\times\mathcal{S}}$ is the instantaneous reward function. The set $\mathcal{U}\subset\left(\Delta(\mathcal{S})\right)^{\mathcal{S}\times\mathcal{A}}$, usually called the uncertainty set, models the set of all possible transition probabilities $\bm{P}=\left(\bm{p}_{sa}\right)_{sa}\in\Delta(\mathcal{S})^{\mathcal{S}\times\mathcal{A}}$, and $\bm{p}_{0}\in\Delta(\mathcal{S})$ is the initial distribution over the set of states $\mathcal{S}$. We emphasize that we assume throughout $\mathcal{S}$ and $\mathcal{A}$ to be finite. We also assume that $\mathcal{U}$ is a non-empty, convex compact set. This is well-motivated from a practical standpoint, where $\mathcal{U}$ is typically built from some data based on some statistical distances, see the end of this section for some examples.
We write $\Pi_{\sf H}$ for the set of history-dependent, randomized policies that possibly depend on the entire past history $(s_{0},a_{0},...,s_{t})$ to choose possibly randomly an action at time $t$. We write $\Pi_{\sf M}$ for the set of Markovian policies, i.e., for policies $\pi\in\Pi_{\sf H}$ that only depend on $t$ and the current state $s_{t}$ for the choice of $a_{t}$. The set $\Pi_{\sf S}\subset\Pi_{M}$ of stationary policies consists of the Markovian policies that do not depend on time, and $\Pi_{\sf SD}\subset\Pi_{\sf S}$ are those policies that are furthermore deterministic (Puterman, 2014). Given a policy $\pi\in\Pi_{\sf H}$ and transition probabilities $\bm{P}\in\mathcal{U}$, we denote by $\mathbb{E}_{\pi,\bm{P}}$ the expectation with respect to the distribution of the sequence $(s_{t},a_{t})_{t\geq 0}$ of states and actions induced by $\pi$ and $\bm{P}$ (as a function of the initial state). Given $\gamma<1$, the value function induced by $(\pi,\bm{P})\in\Pi_{\sf H}\times\mathcal{U}$ is
$v^{\pi,\bm{P}}_{\gamma,s}:=\mathbb{E}_{\pi,\bm{P}}\left[\sum_{t=0}^{+\infty}\gamma^{t}r_{s_{t}a_{t}s_{t+1}}\;|\;s_{0}=s\right],\forall\;s\in\mathcal{S},$
and the discounted return is $R_{\gamma}(\pi,\bm{P}):=\bm{p}_{0}^{\top}\bm{v}^{\pi,\bm{P}}_{\gamma}$. A discount optimal policy is an optimal solution to the optimization problem:
$\sup_{\pi\in\Pi_{\sf H}}\min_{\bm{P}\in\mathcal{U}}\,R_{\gamma}(\pi,\bm{P}).$ (2.1)
Rectangularity, Bellman operator and adversarial MDP. The discounted robust MDP problem (2.1) becomes tractable when the uncertainty set $\mathcal{U}$ satisfies the following s-rectangularity condition:
$\mathcal{U}=\times_{s\in\mathcal{S}}\,\mathcal{U}_{s},\quad\mathcal{U}_{s}\subseteq\Delta(\mathcal{S})^{\mathcal{A}},\quad\forall\;s\in\mathcal{S},$ (2.2)
with the interpretation that $\mathcal{U}_{s}$ is the set of probability transitions from state $s\in\mathcal{S}$, as a function of the action $a\in\mathcal{A}$. The authors in Wiesemann et al. (2013) show that if the uncertainty set $\mathcal{U}$ is s-rectangular and compact convex, a discount optimal policy $\pi^{\star}$ can be chosen stationary (but it may be randomized). Additionally, $\pi^{\star}$ can be computed by first solving for $\bm{v}\in\mathbb{R}^{\mathcal{S}}$ the equation: $\bm{v}^{\star}=T(\bm{v}^{\star})$ where $T:\mathbb{R}^{\mathcal{S}}\rightarrow\mathbb{R}^{\mathcal{S}}$ is the Bellman operator, defined as
$T_{s}(\bm{v})=\max_{\pi\Pi_{S}}\min_{\bm{P}_{s}\in\mathcal{U}_{s}}\;\sum_{a\in\mathcal{A}}\pi_{sa}\bm{p}_{sa}^{\top}\left(\bm{r}_{sa}+\gamma\bm{v}\right),\forall\;\bm{v}\in\mathbb{R}^{\mathcal{S}},\forall\;s\in\mathcal{S},$ (2.3)
then returning the policy $\pi^{\star}$ that attains the maximum on each component of $T(\bm{v}^{\star})$.
Let us now fix a stationary policy $\pi\in\Pi_{\sf S}$, and a s-rectangular RMDP instance. The problem of computing the worst-case return of $\pi$, defined as
$\min_{\bm{P}\in\mathcal{U}}R_{\gamma}(\pi,\bm{P})$ (2.4)
can be reformulated as an MDP instance, called the adversarial MDP. In the adversarial MDP, the state set is the finite set $\mathcal{S}$, the set of actions available at $s\in\mathcal{S}$ is the compact set $\mathcal{U}_{s}$, and the rewards and transitions are continuous, see Ho et al. (2021); Goyal and Grand-Clément (2022) for more details. In particular, the robust value function $\bm{v}^{\pi,\mathcal{U}}_{\gamma}\in\mathbb{R}^{\mathcal{S}}$ of $\pi$, defined as $v^{\pi,\mathcal{U}}_{\gamma,s}:=\min_{\bm{P}\in\mathcal{U}}v^{\pi,\bm{P}}_{\gamma,s}$ for all $s\in\mathcal{S}$, satisfies the fixed-point equation:
$v^{\pi,\mathcal{U}}_{\gamma,s}\;=\;\min_{\bm{P}_{s}\in\mathcal{U}_{s}}\,\sum_{a\in\mathcal{A}}\pi_{sa}\bm{p}_{sa}^{\top}\left(\bm{r}_{sa}+\gamma\bm{v}^{\pi,\mathcal{U}}_{\gamma}\right),\quad\forall\;s\in\mathcal{S}.$
Crucially, the worst-case transition probabilities for a given $\pi\in\Pi_{\sf S}$ can be chosen in the set of extreme points of $\mathcal{U}$.
Under the sa-rectangularity assumption:
$\mathcal{U}=\bigtimes_{(s,a)\in\mathcal{S}\times\mathcal{A}}\,\mathcal{U}_{sa},\quad\mathcal{U}_{sa}\subseteq\Delta(\mathcal{S}),\quad\forall\;(s,a)\in\mathcal{S}\times\mathcal{A},$ (2.5)
a discount optimal policy may be chosen stationary and deterministic when $\mathcal{U}$ is compact (Iyengar, 2005; Nilim and Ghaoui, 2005). Typical examples of sa-rectangular uncertainty sets are based on $\ell_{p}$-distance with $p\in\{1,2,+\infty\}$ , with $\mathcal{U}_{sa}$ defined as $\mathcal{U}_{sa}=\{\bm{p}\in\Delta(\mathcal{S})\;|\;\|\bm{p}-\bm{p}_{sa}\|_{p}\leq\alpha_{sa}\}$ or based on Kullback-Leibler divergence: $\mathcal{U}_{sa}=\{\bm{p}\in\Delta(\mathcal{S})\;|\;\operatorname{KL}\left(\bm{p},\bm{p}_{sa}\right)\leq\alpha_{sa}\}$ for $(s,a)\in\mathcal{S}\times\mathcal{A}$ and some radius $\alpha_{sa}\geq 0$ (Iyengar, 2005; Nilim and Ghaoui, 2005; Panaganti and Kalathil, 2022). Typical s-rectangular uncertainty sets are constructed analogously to sa-rectangular ambiguity sets (Wiesemann et al., 2013; Grand-Clement and Kroer, 2021).
2.2 Connections with stochastic games
Two-player zero-sum stochastic games (abbreviated SGs in the rest of this paper) were introduced in the seminal work of Shapley (Shapley, 1953) and model repeated interactions between two agents with opposite interests. In each period, the players choose each an action, which influences the evolution of the (publicly observed) state. Agents aim at optimizing their gain, which depends on states and actions. The literature on stochastic games is extensive, we refer the reader to Neyman et al. (2003); Mertens et al. (2015) for classical textbooks and to Laraki and Sorin (2015); Renault (2019) for recent short surveys. Stochastic games and robust MDPs share many similarities despite historically distinct communities, motivations, terminology, and lines of research. In this paper, we leverage some existing results in SGs to prove novel results for RMDPs. This section briefly describes the similarities and differences between the two fields.
Similarities between SGs and RMDPs. It has been noted in the RMDP literature that rectangular RMDPs as in (2.1) can be reformulated as stochastic games. At a given state $s\in\mathcal{S}$, the first player in the game is the decision-maker in the RMDP, who chooses actions in $\mathcal{A}$, while the adversary chooses the transition probabilities $\bm{P}_{s}$ in $\mathcal{U}_{s}$. The uncertainty set thus corresponds to the action set of the adversary. This connection has been alluded to multiple times in the robust MDP literature, e.g., section 5 of Iyengar (2005), the last paragraph of page 7 of Xu and Mannor (2010), the last paragraph of section 2 in Wiesemann et al. (2013), and the introduction of Grand-Clément and Petrik (2022).
It is crucial to note that polyhedral uncertainty sets in s-rectangular RMDPs can be modeled by a finite action set for the second player in the associated SG. This is because the worst-case transition probabilities may be chosen in the set of extreme points of $\mathcal{U}$, and polyhedra have finitely many extreme points. With this in mind, extreme points of the s-rectangular convex set $\mathcal{U}$ correspond to stationary deterministic strategies for the second player, while arbitrary elements of $\mathcal{U}$, which are convex combinations of extreme points, correspond to stationary randomized strategies, also called stationary (behavior) strategies in game theory.
Interestingly, sa-rectangular RMDPs can be reformulated as a special case of perfect information stochastic games (see section 5 in Iyengar (2005)). Perfect information SGs are SGs with the property that each state is controlled (in terms of rewards and transitions) by only one player, see Chapter 4 in Neyman et al. (2003) for an introduction. In the perfect information SG reformulation of sa-rectangular RMDPs, the set of states is the union of $\mathcal{S}$ and of $\mathcal{S}\times\mathcal{A}$, the first player controls states $s\in\mathcal{S}$, and the second player chooses the transition probabilities $\bm{p}_{sa}$ in states $(s,a)\in\mathcal{S}\times\mathcal{A}$. Instantaneous rewards are only obtained at the states of the form $(s,a)$. For completeness, a detailed construction is provided in Appendix 8.
A fundamental distinction between SGs and RMDPs. We now describe a crucial distinction between RMDPs and SGs, which has received limited attention in the RMDP literature. In the classical stochastic game framework, the first player and the second player can choose history-dependent strategies. In contrast, in robust MDPs as in (2.1), it is common to assume that the decision-maker chooses a history-dependent policy, but that the adversary is restricted to stationary policies, i.e., to choose $\bm{P}\in\mathcal{U}$ instead of $\bm{P}\in\mathcal{U}_{\sf H}$ with $\mathcal{U}_{\sf H}$ is the set of (randomized) history-dependent policies for the adversary (Wiesemann et al., 2013; Goyal and Grand-Clément, 2022). This fundamental difference between SGs and RMDPs is one of the main reasons why existing results for SGs do not readily extend to results for RMDPs. The focus of this paper is on stationary adversaries, as classical for RMDPs, and for clarity we make it explicit in the statements of all our results.
Remark 2.1
This distinction is irrelevant in the discounted return case, as we point out in Proposition 2.2 below. This proposition follows from Shapley (1953) and we provide a concise proof in Appendix 9. Equality (2.6) shows that facing a stationary adversary or a history-dependent adversary is equivalent for RMDPs with discounted return (and a decision-maker that can choose history-dependent policies). This provides an answer to some of the discussions about stationary vs. non-stationary adversaries from the seminal paper on discounted sa-rectangular RMDPs (Iyengar, 2005).
Proposition 2.2
Let $\mathcal{U}$ be a convex compact s-rectangular uncertainty set. Then
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{\gamma}(\pi,\bm{P})=\sup_{\pi\in\Pi_{\sf H}}\min_{\bm{P}\in\mathcal{U}}R_{\gamma}(\pi,\bm{P}).$ (2.6)
3 Robust MDPs with average optimality
In this section, we study the average optimality criterion for rectangular RMDPs. We start by introducing average optimality, and we discuss the gaps in the existing literature in Section 3.1. We then show in Section 3.2 that stationary deterministic average optimal policies exist for sa-rectangular compact convex uncertainty sets. However and perhaps surprisingly, we show in Section 3.3 for s-rectangular RMDPs that Markovian policies may strictly improve upon stationary ones.
3.1 Average optimality and gaps in the existing literature
Average optimality is a fundamental objective criterion extensively studied in nominal MDPs and in reinforcement learning, see Chapters 8 and 9 in Puterman (2014) and the survey Dewanto et al. (2020). Average optimality alleviates the need for introducing a discount factor by directly optimizing a long-run average $R_{{\sf avg}}(\pi,\bm{P})$ of the instantaneous rewards received over time. Intuitively, we would want $R_{{\sf avg}}(\pi,\bm{P})$ to capture the limit behaviour of the average payoff $\mathbb{E}_{\pi,\bm{P}}\left[\frac{1}{T+1}\sum_{t=0}^{T}r_{s_{t}a_{t}s_{t+1}}\;|\;s_{0}\sim\bm{p}_{0}\right]$ over the first $T$ periods, for $T\in\mathbb{N}$. A well-known issue however, see for instance Example 8.1.1 in Puterman (2014), is that these average payoffs need not have a limit as $T\rightarrow+\infty$.
Example 3.1
Consider a MDP with a single state and two actions $a_{0}$ and $a_{1}$, with payoffs 0 and 1 respectively. In this setup, a (deterministic) Markovian strategy is a sequence $(a_{t})$ in $\{0,1\}^{\mathbb{N}}$, and the average payoff over the first $T$ periods is the frequency of action $a_{1}$ in these periods. Therefore, if the sequence $(a_{t})_{t\in\mathbb{N}}$ is such that these frequencies do not converge, the average payoff up to $T$ does not converge either as $T\to+\infty$. We provide a detailed example in Appendix 10.
In this paper, we focus on the following definition of the average return $R_{{\sf avg}}(\pi,\bm{P})$ of a pair $\left(\pi,\bm{P}\right)\in\Pi_{\sf H}\times\mathcal{U}$:
$R_{{\sf avg}}(\pi,\bm{P})=\mathbb{E}_{\pi,\bm{P}}\left[\limsup_{T\rightarrow+\infty}\frac{1}{T+1}\sum_{t=0}^{T}r_{s_{t}a_{t}s_{t+1}}\;|\;s_{0}\sim\bm{p}_{0}\right].$ (3.1)
[…]
$R_{{\sf avg}}(\pi,\bm{P})=\lim_{T\rightarrow+\infty}\mathbb{E}_{\pi,\bm{P}}\left[\frac{1}{T+1}\sum_{t=0}^{T}r_{s_{t}a_{t}s_{t+1}}\;|\;s_{0}\sim\bm{p}_{0}\right],\forall\;(\pi,\bm{P})\in\Pi_{\sf S}\times\mathcal{U},$
and the average return coincides with the limit (normalized) discounted return, as the discount factor increases to 1:
$\lim_{\gamma\rightarrow 1}(1-\gamma)R_{\gamma}(\pi,\bm{P})=R_{{\sf avg}}(\pi,\bm{P}),\forall\;(\pi,\bm{P})\in\Pi_{\sf S}\times\mathcal{U}.$
One of our main goals in this section is to study the robust MDP problem with the average return criterion:
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$ (3.2)
A solution $\pi\in\Pi_{\sf H}$ to (3.2), when it exists, is called an average optimal policy. The definition of the optimization problem (3.2) warrants two important comments.
Properties of average optimal policies. First, we stress that (3.2) optimizes the worst-case average return over history-dependent policies. In contrast, prior work on average optimal RMDPs (Tewari and Bartlett, 2007; Le Tallec, 2007; Wang et al., 2023) have solely considered the optimization problem $\sup_{\pi\in\Pi_{\sf S}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$, thus restricting the decision-maker to stationary policies. At this point, it is not clear if stationary policies are optimal in (3.2), and this issue has been overlooked in the existing literature. One of our main contributions is to answer this question by the positive for sa-rectangular RMDPs and by the negative for s-rectangular RMDPs.
Non-existence of the worst-case transition probabilities. Second, we note that (3.2) considers the infimum over $\bm{P}\in\mathcal{U}$ of the average return $R_{{\sf avg}}(\pi,\bm{P})$. The reason is that this infimum may not be attained, even for $\pi\in\Pi_{\sf S}$. This issue has been overlooked in prior work on robust MDPs with average return (Tewari and Bartlett, 2007; Wang et al., 2023), even though it was recognized in other communities repeatedly, see e.g. Section 1.4.4 in Sorin (2002) or Example 4.10 in Leizarowitz (2003). In the next proposition, we provide a simple counterexample.
Proposition 3.2
There exists a robust MDP instance with an sa-rectangular compact convex uncertainty set, for which $\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$ is not attained for any $\pi\in\Pi_{\sf S}$.
The proof builds upon a relatively simple RMDP instance with three states and one action. We also reuse this RMDP instance for later important results in the next sections. For this reason, we describe it in detail below.
Proof.
Proof of Proposition 3.2. We consider the robust MDP instance from Figure 1.
(a) The point $(1-\alpha,\beta,\alpha-\beta)$ in $\Delta(\{s_{0},s_{1},s_{2}\})$
(b) The uncertainty set $\mathcal{U}_{s_{0}a_{0}}$ (in blue)
Figure 1: A simple robust MDP instance where $\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$ is not attained.
There are three states: $\mathcal{S}=\{s_{0},s_{1},s_{2}\}$, and only one action $a_{0}$, so that the RMDP reduces to the adversarial MDP with the adversary aiming at minimizing the average return. States $s_{1}$ and $s_{2}$ are absorbing states with reward $-1$ and 00 respectively. The initial state is $s_{0}$, and the instantaneous reward in $s_{0}$ is 00. Elements of $\Delta(\mathcal{S})$ are written $(p_{0},p_{1},p_{2})$, where $p_{i}$ is the probability of transitioning to state $s_{i}$. The uncertainty set is
$\mathcal{U}_{s_{0}a_{0}}=\{\left(1-\alpha,\beta,\alpha-\beta\right)\in\Delta(\mathcal{S})\;|\;\alpha,\beta\in[0,1],\beta\leq\alpha(1-\alpha)\},$
see Figure 0(a). Given $\left(1-\alpha,\beta,\alpha-\beta\right)\in\mathcal{U}_{s_{0}a_{0}}$, $\alpha$ is the probability to leave state $s_{0}$ and $\beta$ as the probability to go to state $s_{1}$.
The decision-maker has a single policy, which we denote $\pi=a_{0}$. We claim that $\inf_{\bm{P}\in\mathcal{U}_{s_{0}a_{0}}}R_{{\sf avg}}(\pi,\bm{P})=-1$, with $R_{{\sf avg}}(\pi,\bm{P})>-1$ for each $\bm{P}\in\mathcal{U}_{s_{0}a_{0}}$. We first compute the average return associated with a pair $(a_{0},\bm{P})$ with $\bm{P}\in\mathcal{U}_{s_{0}a_{0}}$, $\bm{P}=(1-\alpha,\beta,\alpha-\beta)$ with $\alpha,\beta\in[0,1]$. Given the instantaneous rewards in this instance, note that we always have $R_{{\sf avg}}(\pi,\bm{P})\in[-1,0]$. If $\alpha=0$, then the decision-maker never leaves $s_{0}$ and $R_{{\sf avg}}(\pi,\bm{P})=0.$ Otherwise, for every discount factor $\gamma\in(0,1)$, the discounted return $R_{\gamma}(a_{0},\bm{P})$ satisfies
$(1-\gamma)R_{\gamma}(a_{0},\bm{P})=-\gamma\frac{\beta}{1-\gamma+\gamma\alpha}.$
[…]
Prior works on RMDPs have only focused on sa-rectangular RMDPs, either with polyhedral uncertainty set (Le Tallec, 2007; Tewari and Bartlett, 2007) or under the assumption that the Markov chains induced by any pair of policy and transition probabilities are unichain (Wang et al., 2023). In these special cases, we prove in Appendix 11 that $\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})=\min_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$, i.e., worst-case transition probabilities indeed exist for any stationary policy $\pi\in\Pi_{\sf S}$. We note that this question is not mentioned in prior work (Tewari and Bartlett, 2007; Wang et al., 2023).
We conclude this section with the following technical lemma. We adapt it from the literature on nominal MDP with a finite set of states and compact set of actions (Bierth, 1987). We briefly discuss this lemma in Appendix 12.
Lemma 3.3 (Adapted from theorem 2.5, Bierth (1987))
Let $\mathcal{U}$ be compact convex and s-rectangular. Let $\pi\in\Pi_{\sf S}$. Then
$\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})=\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P}).$
3.2 The case of sa-rectangular robust MDPs
We now focus on the case of sa-rectangular RMDPs. Our main result in this section is to show that there always exist average optimal policies that are stationary and deterministic, an attractive feature for practical implementation in real-world applications. In particular, we have the following theorem.
Theorem 3.4
Consider a sa-rectangular robust MDP with a compact convex uncertainty set $\mathcal{U}$. There exists an average optimal policy that is stationary and deterministic:
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})\;=\;\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P}).$
The proof proceeds in two steps and leverages existing results from the literature on perfect information SGs (Gimbert and Kelmendi, 2023).
Proof.
Proof of Theorem 3.4. In the first step of the proof, we show that Theorem 3.4 is true in the special case where $\mathcal{U}$ is polyhedral. In this case, $\mathcal{U}$ only has a finite number of extreme points, and the robust MDP problem is equivalent to a perfect information stochastic game with finitely many actions for both players, as described in Section 2. We can then rely on the following crucial equality, which is a reformulation of proposition 3.2 in Gimbert and Kelmendi (2023):
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})=\inf_{\bm{P}\in\mathcal{U}}\sup_{\pi\in\Pi_{\sf H}}R_{{\sf avg}}(\pi,\bm{P}).$ (3.3)
From weak duality we always have $\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})\leq\inf_{\bm{P}\in\mathcal{U}}\sup_{\pi\in\Pi_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})$, and
$\inf_{\bm{P}\in\mathcal{U}}\sup_{\pi\in\Pi_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})\leq\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})\leq\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$
where the equality follows from (3.3), the second inequality follows from $\mathcal{U}\subset\mathcal{U}_{\sf H}$, and the last inequality follows from $\Pi_{\sf SD}\subset\Pi_{\sf H}$. Therefore, all terms above are equal, and we have
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P}).$
In the second step of the proof, we show that Theorem 3.4 holds for general compact uncertainty sets, without the assumption that $\mathcal{U}$ is polyhedral as in the first step of the proof. Let $\epsilon>0.$ For each $\pi\in\Pi_{\sf SD}$, consider $\bm{P}^{\pi,\epsilon}\in\mathcal{U}$ such that
$\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})\leq R_{{\sf avg}}(\pi,\bm{P}^{\pi,\epsilon})\leq\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})+\epsilon.$
Note that $\bm{P}^{\pi,\epsilon}$ always exists by definition of the infimum and note that $\{\bm{P}^{\pi,\epsilon}\;|\;\pi\in\Pi_{\sf SD}\}$ is a finite set since $\Pi_{\sf SD}$ is a finite set. Let us now consider the sa-rectangular uncertainty set $\mathcal{U}_{\sf f}$ defined as the convex hull of the finite set $\{\bm{P}\;|\;\forall\;(s,a)\in\mathcal{S}\times\mathcal{A},\exists\pi\in\Pi_{\sf SD},\bm{p}_{sa}=\bm{p}_{sa}^{\pi}\}.$ By construction, $\mathcal{U}_{\sf f}$ is polyhedral and sa-rectangular. Additionally, $\mathcal{U}_{\sf f}\subset\mathcal{U}$, and for any $\pi\in\Pi_{\sf SD}$, we have
$\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})\leq\inf_{\bm{P}\in\mathcal{U}_{\sf f}}R_{{\sf avg}}(\pi,\bm{P})\leq\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})+\epsilon.$
Therefore,
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$ $\leq\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}_{\sf f}}R_{{\sf avg}}(\pi,\bm{P})$
$=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}_{\sf f}}R_{{\sf avg}}(\pi,\bm{P})$
$\leq\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})+\epsilon.$
where the first inequality uses $\mathcal{U}_{\sf f}\subset\mathcal{U}$, the equality follows from the first step of the proof and $\mathcal{U}_{\sf f}$ being polyhedral, and the last equality holds by construction of $\mathcal{U}_{\sf f}$. Therefore, for all $\epsilon>0$, we have $\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})\leq\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})+\epsilon$, and we can conclude that
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})\leq\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$
which directly implies
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P}).$
[…]
Theorem 3.4 has several important consequences.
Closing an important gap in prior works. First, to the best of our knowledge, we are the first to study average return RMDPs in all generality without constraining the problem to stationary policies and to show that average optimal policies for sa-rectangular RMDPs may be chosen stationary and deterministic. Therefore, Theorem 3.4 addresses an important gap that has been entirely overlooked in the existing literature on RMDPs with average return (Tewari and Bartlett, 2007; Wang et al., 2023).
Strong duality and history-dependent adversary. Another consequence of Theorem 3.4 is the following strong duality theorem. We provide the detailed proof in Appendix 13.
Theorem 3.5
Consider a sa-rectangular robust MDP with a compact convex uncertainty set $\mathcal{U}$. Then the following strong duality results hold:
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$ $=\inf_{\bm{P}\in\mathcal{U}}\sup_{\pi\in\Pi_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})$ (3.4)
$\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$ $=\inf_{\bm{P}\in\mathcal{U}}\max_{\pi\in\Pi_{\sf SD}}R_{{\sf avg}}(\pi,\bm{P})$ (3.5)
Theorem 3.5 is akin to the strong duality results for discounted RMDPs (Wiesemann et al., 2013; Goyal and Grand-Clément, 2022). It is interesting to note that strong duality still holds for sa-rectangular RMDPs with average optimality, especially as in Equality (3.5), since the maximum over $\pi\in\Pi_{\sf SD}$ is always attained (it is the maximum over a finite set) while the infimum over $\mathcal{U}$ may not be attained, even in very simple settings, as we illustrate in Proposition 3.2. Strong duality is crucial to study the case of history-dependent adversaries, as we now show. The case of non-stationary adversary has also gathered interest and is discussed in Iyengar (2005); Nilim and Ghaoui (2005). Interestingly, we show that this model is equivalent to (3.9), i.e., to the case of stationary adversaries in the following theorem. The proof is relatively concise and relies on our fundamental results from Theorem 3.4, Theorem 3.5 and on the properties of MDPs with compact action sets. We also note that it is straightforward to prove that the same results hold for Markovian adversaries.
Theorem 3.6
Consider a sa-rectangular robust MDP with a compact convex uncertainty set $\mathcal{U}$. Then
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})\;=\;\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P}).$
Proof.
Proof. We have
$\inf_{\bm{P}\in\mathcal{U}_{\sf H}}\sup_{\pi\in\Pi_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})$ $\leq\inf_{\bm{P}\in\mathcal{U}}\sup_{\pi\in\Pi_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})$
$=\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$
$=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})$
$=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})$
$\leq\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})$
[…]
Other natural definitions of average optimality. Other possible definitions of the average return criterion exist, such as, for $(\pi,\bm{P})\in\Pi_{\sf H}\times\mathcal{U}$, the following natural definitions:
$\liminf_{T\rightarrow+\infty}\mathbb{E}_{\pi,\bm{P}}\left[\frac{1}{T+1}\sum_{t=0}^{T}r_{s_{t}a_{t}s_{t+1}}\;|\;s_{0}\sim\bm{p}_{0}\right],$ (3.6)
$\limsup_{T\rightarrow+\infty}\mathbb{E}_{\pi,\bm{P}}\left[\frac{1}{T+1}\sum_{t=0}^{T}r_{s_{t}a_{t}s_{t+1}}\;|\;s_{0}\sim\bm{p}_{0}\right],$ (3.7)
$\mathbb{E}_{\pi,\bm{P}}\left[\liminf_{T\rightarrow+\infty}\frac{1}{T+1}\sum_{t=0}^{T}r_{s_{t}a_{t}s_{t+1}}\;|\;s_{0}\sim\bm{p}_{0}\right].$ (3.8)
Note that when $(\pi,\bm{P})\in\Pi_{\sf S}\times\mathcal{U}$, the three quantities in the above equations are equal to $R_{{\sf avg}}$ and are equal to
$\lim_{T\rightarrow+\infty}\;\mathbb{E}_{\pi,\bm{P}}\left[\frac{1}{T+1}\sum_{t=0}^{T}r_{s_{t}a_{t}s_{t+1}}\;|\;s_{0}\sim\bm{p}_{0}\right]$ (3.9)
but in all generality, we have the following inequality (see lemma 2.1 in Feinberg and Shwartz (2012)):
[…]
Corollary 3.7
Consider a sa-rectangular robust MDP with a compact convex uncertainty set $\mathcal{U}$. Let $\hat{R}_{{\sf avg}}$ as defined in (3.6), (3.7), or (3.8).
  1. 1.
    There exists an average optimal policy that is stationary and deterministic, and which coincides with an average optimal policy for $R_{{\sf avg}}$ as in (3.1):
    $\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}}\hat{R}_{{\sf avg}}(\pi,\bm{P})\;=\;\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}\hat{R}_{{\sf avg}}(\pi,\bm{P})=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P}).$
  2. 2.
    The strong duality results from Theorem 3.5 still hold if we replace $R_{{\sf avg}}$ by $\hat{R}_{{\sf avg}}$.
  3. 3.
    The equivalence between stationary adversaries and history-dependent adversaries from Theorem 3.6 still hold if we replace $R_{{\sf avg}}$ by $\hat{R}_{{\sf avg}}$.
3.3 The case of s-rectangular robust MDPs
We now show that for s-rectangular RMDPs, Markovian policies may strictly outperform stationary policies for the average return criterion. This is a surprising result since stationary policies are sufficient for solving average sa-rectangular RMDPs and discount s-rectangular RMDPs. We rely on a stochastic game instance called The Big Match (Blackwell and Ferguson, 1968). In this instance, there are two absorbing states $A_{0}$ (with a reward of 00) and $A_{1}$ (with a reward of $1$) and two non-absorbing state $s_{0}$ and $s_{1}$. The decision-maker starts in state $s_{0}$ and chooses an action in $\{T,B\}$. The adversary chooses the scalar $p\in[0,1]$, whose impact on the transitions is represented in Figure 2. Each arrow corresponds to a possible transition, and is labelled with its probability, and with the payoff obtained along this transition. If action $T$ is chosen, the game reaches an absorbing state, $A_{0}$ or $A_{1}$, depending on the choice of the adversary. If action $B$ is chosen, the decision-maker obtains a reward of 00 or $1$ and the game continues. We can think of this instance as a game where the adversary tries to “guess” if the decision-maker will choose to stop the game (action $T$). If the adversary guesses correctly and the decision-maker chooses action $T$, the adversary can choose $p=1$ so that the decision-maker reaches the “bad” absorbing state $A_{0}$ (with reward 00). Otherwise, if the decision-maker chooses action $B$, the game continues. The Big Match is represented in Figure 2. We will show that in The Big Match, all stationary policies have a worst-case average return of 00, while a Markovian policy can achieve a worst-case average return of $1/2$. In particular, we have the following proposition.
(a) Transition for action $T$
(b) Transition for action $B$
Figure 2: Transitions and rewards for The Big Match (Blackwell and Ferguson, 1968; Gillette, 1957) reformulated as an s-rectangular RMDP. The adversary chooses $p\in[0,1]$ and the decision-maker chooses an action in $\{T,B\}$.
Proposition 3.8
Consider the robust MDP instance from Figure 2 (The Big Match).
  1. 1.
    For any history-dependent policy $\pi\in\Pi_{\sf H}$, we have $\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})\leq 1/2$.
  2. 2.
    There exists a Markovian policy $\pi\in\Pi_{\sf M}$ such that $\min_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})=1/2$.
  3. 3.
    For any stationary policy $\pi\in\Pi_{\sf S}$, we have $\min_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})=0$. Additionally, $\min_{\bm{P}\in\mathcal{U}}\max_{\pi\in\Pi_{\sf S}}R_{{\sf avg}}(\pi,\bm{P})=1/2$. Therefore, there is a duality gap in the case where the decision-maker is constrained to choose a stationary policy:
    $\max_{\pi\in\Pi_{\sf S}}\min_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})=0<\min_{\bm{P}\in\mathcal{U}}\max_{\pi\in\Pi_{\sf S}}R_{{\sf avg}}(\pi,\bm{P})=1/2.$
Proof.
Proof. We note that point 1 and point 3 are known in the SG literature, and only point 2 is new here. In particular, point 1 follows from Blackwell’s original argument, see the beginning of theorem 1 in Blackwell and Ferguson (1968), and point 3 is exactly Lemma 1, chapter 12, Neyman et al. (2003). We detail the proofs here for conciseness.
  1. 1.
    This follows from the adversary choosing the transition probabilities $\bm{P}$ corresponding to $p=1/2$. In this case, we have $R_{{\sf avg}}(\pi,\bm{P})=1/2$ for any policy $\pi\in\Pi_{\sf H}$. Indeed, if $\pi$ always selects action $B$, then clearly $R_{{\sf avg}}(\pi,\bm{P})=1/2$. Otherwise, there exists a period at which $\pi$ selects action $T$ with a positive probability, and in this case we also have $R_{{\sf avg}}(\pi,\bm{P})=1/2$ since the decision-maker is equally likely to reach the terminal state $A_{0}$ (with a reward of 00) or the terminal state $A_{1}$ (with a reward of $1$).
  2. 2.
    We will construct a Markovian policy $\pi$ such that $R_{{\sf avg}}(\pi,\bm{P})=1/2$ for any $\bm{P}\in\mathcal{U}$. In particular, let $\pi$ the Markovian policy that chooses the same action in the states $s_{0}$ and $s_{1}$ and such that: (a) at time $t=0$, $\pi$ chooses action $T$ or action $B$ with probability $1/2$; (b) for any time $t\geq 1$, $\pi$ always chooses action $B$. Let $p\in[0,1]$ and let $\bm{P}\in\mathcal{U}$ be the corresponding transition probabilities. At time $t=0$, the game stops with probability $1/2$ (if action $T$ is chosen), in which case the decision-maker reaches $A_{0}$ with probability $p$, obtaining a reward of 00 forever, and reaches $A_{1}$ with probability $1-p$, obtaining a reward of $1$ forever. If action $B$ is selected at time $t=0$, which happens with probability $1/2$, the game starts in $s_{0}$ or $s_{1}$ at time $t=1$. Then, action $B$ is selected forever, so that the average payoff is $(1-p).$ Overall, we obtain that the average payoff of $\pi$ is $\frac{1}{2}\times p+\frac{1}{2}\times(1-p)=\frac{1}{2}$.
  3. 3.
    Let $\pi\in\Pi_{\sf S}$. If $\pi$ always chooses action $B$, then the adversary choosing $p=0$ yields an average return of 00. Otherwise, $\pi$ chooses action $T$ with a positive probability, and the adversary choosing $p=1$ yields an average return of 00. Therefore, $\min_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P})=0$ for any $\pi\in\Pi_{\sf S}$. The fact that $\min_{\bm{P}\in\mathcal{U}}\max_{\pi\in\Pi_{\sf S}}R_{{\sf avg}}(\pi,\bm{P})=1/2$ follows from point 1 in this proof: when $p=1/2$, we have $R_{{\sf avg}}(\pi,\bm{P})=1/2,\forall\;\pi\in\Pi_{\sf H}$.
[…]
Proposition 3.8 shows the stark contrast between average optimal policies for sa-rectangular RMDPs, which can always be chosen stationary and deterministic (Theorem 3.4), and average optimal policies for s-rectangular RMDPs, which may have to be Markovian and randomized. Additionally, the last point of Proposition 3.8 shows that strong duality does not hold for average return s-rectangular RMDPs with stationary strategies for the decision-maker, again in contrast to the case of sa-rectangular RMDPs (Theorem 3.5). It is an interesting open question to understand if, for any s-rectangular RMDP, an average optimal policy may be chosen Markovian and if strong duality holds in the case of Markovian decisions for the decision-maker. We conclude this section with a comparison of The Big Match in the case of a history-dependent adversary. In this case, The Big Match is exactly the instance studied in the stochastic game literature. In particular, Blackwell and Ferguson (1968) show that in the classical SG setting where both the decision-maker and the adversary can use history-dependent policies, we have
$\sup_{\pi\in\Pi_{\sf H}}\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})=\inf_{\bm{P}\in\mathcal{U}_{\sf H}}\sup_{\pi\in\Pi_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})=1/2,$
the supremum is not attained in the left-hand side above, any Markovian policy achieves a worst-case average return of 00: $\inf_{\bm{P}\in\mathcal{U}_{\sf H}}R_{{\sf avg}}(\pi,\bm{P})=0$ for $\pi\in\Pi_{\sf M}$, and for any choice of $\epsilon\in(0,1/2)$, only history-dependent policies can guarantee a worst-case average reward of $1/2-\epsilon$. We refer to chapter 12 in Neyman et al. (2003) for a modern exposition of these results.
4 Robust MDPs with Blackwell optimality
We now study Blackwell optimality for rectangular RMDPs. As detailed in the introduction, Blackwell optimality provides an adequate optimality criterion when there is no natural notion of discounting. We first introduce (approximate) Blackwell optimality in Section 4.1. We then focus on sa-rectangular RMDPs in Section 4.2 and on s-rectangular RMDPs in Section 4.3.
4.1 Blackwell optimality and approximate Blackwell optimality
An important limitation of the average return is that it ignores any rewards obtained in finite time, which may be problematic. For instance, when optimizing patient trajectories over time, practitioners are concerned with long-term goals (typically, survival at discharge) but also with the current patient condition at any point in time. Blackwell optimality is a criterion that balances both long-term and short-term goals by accounting for an entire range of discount factors. In particular, a policy is Blackwell optimal if it is discount optimal for all discount factors sufficiently close to $1$ (Puterman, 2014).
Definition 4.1
Let $\mathcal{U}$ be s-rectangular and convex compact. A policy $\pi\in\Pi_{\sf H}$ is Blackwell optimal if there exists $\gamma_{0}\in(0,1)$ such that
$\min_{\bm{P}\in\mathcal{U}}R_{\gamma}(\pi,\bm{P})\geq\sup_{\pi\in\Pi_{\sf S}}\min_{\bm{P}\in\mathcal{U}}R_{\gamma}(\pi^{\prime},\bm{P}),\quad\forall\;\gamma\in(\gamma_{0},1).$ (4.1)
For nominal MDPs, i.e. for the case where $\mathcal{U}$ is a singleton and the sets $\mathcal{S},\mathcal{A}$ are finite, there always exists a Blackwell optimal policy (Puterman, 2014). We will be interested in the existence (or not) of Blackwell optimal policies for rectangular robust MDPs. We also introduce the following definition of approximate Blackwell optimality, where a policy remains $\epsilon$-optimal for all discount factors sufficiently large.
Definition 4.2 ($\epsilon$-Blackwell optimality.)
Let $\mathcal{U}$ be s-rectangular and convex compact. A policy $\pi\in\Pi_{\sf H}$ is $\epsilon$-Blackwell optimal for $\epsilon>0$ if there exists a discount factor $\gamma_{\epsilon}\in(0,1)$ such that
$\min_{\bm{P}\in\mathcal{U}}\,(1-\gamma)R_{\gamma}(\pi,\bm{P})\;\geq\;\sup_{\pi\in\Pi_{\sf S}}\min_{\bm{P}\in\mathcal{U}}\,(1-\gamma)R_{\gamma}(\pi^{\prime},\bm{P})-\epsilon,\quad\forall\;\gamma\in(\gamma_{\epsilon},1).$ (4.2)
Note that we renormalize the discounted returns in the robust MDPs with the multiplicative term $(1-\gamma)$ since they may be unbounded otherwise. To the best of our knowledge, we are the first to introduce and study the notion of $\epsilon$-Blackwell optimality for robust MDPs. In the rest of this section, we will repeatedly use the following result pertaining to the existence of approximate Blackwell optimal policy in the adversarial MDP, which we adapt from the literature on MDPs with compact action sets.
Theorem 4.3 (Corollary 5.26, Sorin (2002))
Let $\mathcal{U}$ be a compact convex s-rectangular uncertainty set. Let $\pi\in\Pi_{\sf S}$ and $\epsilon>0$. Then there exist $\gamma_{\epsilon}\in(0,1)$ and $\bm{P}_{\epsilon}\in\mathcal{U}$ such that
$\min_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi,\bm{P})\leq(1-\gamma)R_{\gamma}(\pi,\bm{P}_{\epsilon})\leq\min_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi,\bm{P})+\epsilon,\forall\;\gamma\in(\gamma_{\epsilon},1).$
4.2 The case of sa-rectangular robust MDPs
In this section, we provide a complete analysis of Blackwell optimality for sa-rectangular RMDPs. In Section 4.2.1, we first show that, surprisingly, a Blackwell optimal policy may not exist, although an approximate Blackwell optimal policy may always be chosen stationary and deterministic. Additionally, this approximate Blackwell optimal policy is average optimal, as we show in Section 4.2.2. Finally, in Section 4.2.3 we introduce the notion of definable uncertainty sets, a very general class of uncertainty for which a Blackwell optimal policy exists.
4.2.1 Existence and non-existence results
We first contrast the existence properties of Blackwell optimal policies and $\epsilon$-Blackwell optimal policies (as introduced in Definition 4.2). Surprisingly, there are some examples of sa-rectangular robust MDPs where there are no Blackwell optimal policies, as stated formally in the next theorem.
Theorem 4.4
There exists a sa-rectangular robust MDP instance, with a compact convex uncertainty set $\mathcal{U}$, such that there is no Blackwell optimal policy:
$\forall\;\pi\in\Pi_{\sf H},\forall\;\gamma\in(0,1),\exists\;\gamma^{\prime}\in(\gamma,1),\min_{\bm{P}\in\mathcal{U}}R_{\gamma^{\prime}}(\pi,\bm{P})<\sup_{\pi^{\prime}\in\Pi_{\sf S}}\min_{\bm{P}\in\mathcal{U}}R_{\gamma^{\prime}}(\pi^{\prime},\bm{P}).$
To the best of our knowledge, we are the first to show that Blackwell optimal policies may not exist for sa-rectangular RMDPs. Indeed, Theorem 4.4 is surprising because the existence of Blackwell optimal policies has been proved in various other related frameworks, including nominal MDPs (Puterman, 2014), or sa-rectangular RMDPs under some additional assumptions (Goyal and Grand-Clément, 2022; Wang et al., 2023). For the sake of conciseness, we defer the proof of Theorem 4.4 in Appendix 15, and we only provide here some intuition on the main reasons behind the potential non-existence of Blackwell optimal policies.
Our proof of Theorem 4.4 is based on the same simple counterexample as for Proposition 3.2. We consider the same instance but where there are two available actions $a_{1}$ and $a_{2}$ at the non-absorbing state $s_{0}$. We can identify stationary policies with actions $a_{1}$ and $a_{2}$. The main property of our counterexample is that the robust value functions $\gamma\mapsto v^{a_{1},\mathcal{U}}_{s_{0},\gamma}$ and $\gamma\mapsto v^{a_{2},\mathcal{U}}_{s_{0},\gamma}$ have an oscillatory behaviour when $\gamma$ approaches $1$. As these robust value functions oscillate more and more often as $\gamma\rightarrow 1$, they intersect infinitely often on any interval close to $1$, so that there are no discount factors close enough to $1$ after which $a_{1}$ (or $a_{2}$) always remains a discount optimal policy. To obtain the oscillatory behaviours of the robust value functions, we construct two convex compact uncertainty sets $\mathcal{U}_{s_{0}a_{1}}$ and $\mathcal{U}_{s_{0}a_{2}}$. At a high level, we construct these uncertainty sets so that their boundaries overlap and intersect infinitely often, while remaining distinct sets (as subsets of $\Delta(\mathcal{S})$). Let us define the worst-case transition probabilities $\gamma\mapsto\bm{P}^{a_{1},\star}_{\gamma}$ and $\gamma\mapsto\bm{P}^{a_{2},\star}_{\gamma}$ for actions $a_{1}$ and $a_{2}$. Then these worst-case transition probabilities vary with $\gamma$ and always belong to the boundaries of $\mathcal{U}_{s_{0}a_{1}}$ and $\mathcal{U}_{s_{0}a_{2}}$. The infinite intersections of these boundaries induce a pathological, oscillating behaviour of $\gamma\mapsto v^{a_{1},\mathcal{U}}_{s_{0},\gamma}=v^{a_{1},\bm{P}^{a_{1},\star}_{\gamma}}_{s_{0},\gamma}$ (and similarly for $a_{2}$). Our analysis also explains why this pathological behaviour is impossible in nominal MDPs: the transition probabilities is fixed, and in this case $\gamma\mapsto v^{\pi,\bm{P}}_{s,\gamma}$ is always a well-behaved (rational) function. This pathological behaviour is also impossible for polyhedral uncertainty: the worst-case transition probabilities remain the same for $\gamma$ large enough, as shown in (Goyal and Grand-Clément, 2022).
Since Theorem 4.4 shows that Blackwell optimal policies may not always exist, it is natural to ask if approximate Blackwell optimal policies always exist. We answer this question by the positive in the next theorem. In fact, we show the following stronger existence result.
Theorem 4.5
Let $\mathcal{U}$ be a sa-rectangular compact uncertainty set. Then there exists a stationary deterministic policy that is $\epsilon$-Blackwell optimal for all $\epsilon>0$, i.e., $\exists\;\pi\in\Pi_{\sf SD},\forall\;\epsilon>0,\exists\;\gamma_{\epsilon}\in(0,1)$ such that
$\min_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi,\bm{P})\geq\sup_{\pi^{\prime}\in\Pi_{\sf S}}\min_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi^{\prime},\bm{P})-\epsilon,\forall\;\gamma\in(\gamma_{\epsilon},1).$
Theorem 4.5 shows another surprising result: not only there exists an $\epsilon$-Blackwell optimal policy for any choice of $\epsilon>0$, but in fact, we can choose the same stationary deterministic policy $\pi\in\Pi_{\sf SD}$ to be $\epsilon$-Blackwell optimal for all $\epsilon.$ Theorem 4.4 shows that we can not choose $\epsilon=0$ in the statement of Theorem 4.5. We present the proof in Appendix 16. The main idea is to exploit Theorem 4.3, i.e., the existence of approximate Blackwell optimal policies in the adversarial MDP.
4.2.2 Limit behaviours and connection with average optimality
We now highlight the connection between Blackwell optimality and average optimality. Intuitively, approximate Blackwell optimal policies remain $\epsilon$-optimal for all $\gamma$ close to $1$. Since the discount factor captures the willingness of the decision-maker to wait for future rewards, we expect that the discounted return resembles more and more the average return as $\gamma\rightarrow 1$. We show that this intuition is correct for sa-rectangular uncertainty sets in the following theorem.
Theorem 4.6
Consider a sa-rectangular robust MDP with a compact convex uncertainty set $\mathcal{U}$. Let $\pi\in\Pi_{\sf SD}$ be $\epsilon$-Blackwell optimal, for all $\epsilon>0$. Then $\pi$ is also average optimal.
The proof of Theorem 4.6 is presented in Appendix 17. The main lines of the proof are instructive and rely on carefully inspecting the limit behavior of the discounted return as $\gamma\rightarrow 1$. We provide an outline here. Recall that we always have
$\lim_{\gamma\rightarrow 1}(1-\gamma)R_{\gamma}(\pi,\bm{P})=R_{{\sf avg}}(\pi,\bm{P}),\forall\;(\pi,\bm{P})\in\Pi_{\sf S}\times\mathcal{U}.$
The first step is to show that the equality above is still true when taking the worst-case over the transition probabilities. In particular, we show the following lemma.
Lemma 4.7
In the setting of Theorem 4.6, let $\pi\in\Pi_{\sf S}.$ Then $\min_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi,\bm{P})$ admits a limit as $\gamma\rightarrow 1$ and
$\lim_{\gamma\rightarrow 1}\min_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi,\bm{P})=\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P}).$
We then show that the same conclusion holds for the limit of the optimal discounted return.
Lemma 4.8
In the setting of Theorem 4.6, $\max_{\pi\in\Pi_{\sf SD}}\min_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi,\bm{P})$ admits a limit as $\gamma\rightarrow 1$ and
$\lim_{\gamma\rightarrow 1}\max_{\pi\in\Pi_{\sf SD}}\min_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi,\bm{P})=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}R_{{\sf avg}}(\pi,\bm{P}).$
The last part of the proof relates average optimality and approximate-Blackwell optimality.
Lemma 4.9
In the setting of Theorem 4.6, let $\pi\in\Pi_{\sf SD}$ be $\epsilon$-Blackwell optimal for any $\epsilon>0$. Then
$\lim_{\gamma\rightarrow 1}\min_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi,\bm{P})=\lim_{\gamma\rightarrow 1}\max_{\pi^{\prime}\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}(1-\gamma)R_{\gamma}(\pi^{\prime},\bm{P}).$
[…]
Remark 4.10
Lemma 4.8 and Lemma 4.9 are present as corollary 4 in Tewari and Bartlett (2007), under the assumption that $\mathcal{U}$ is based on $\ell_{\infty}$-distance. This considerably simplifies the proof, as in this setting, the number of extreme points of $\mathcal{U}$ is finite, so the infimum over $\mathcal{U}$ can be reduced to a minimum over a finite number of elements, and we can exchange the minimization with the limit. Lemma 4.7 is also present in Wang et al. (2023), under some additional assumptions (unichain compact sa-rectangular RMDPs). Our results show that the unichain assumption is unnecessary.
4.2.3 Definable robust Markov decision processes
In this section, we introduce a very general class of uncertainty sets that encompasses virtually all the practical examples existing in the RMDP literature, and for which stationary deterministic Blackwell optimal policies exist. Indeed, it is classical to construct uncertainty sets based on simple functions, like affine maps, $\ell_{p}$-balls, or Kullback-Leibler divergence, see Section 2. For this kind of simple functions, intuitively we do not expect that the robust value functions oscillate and intersect infinitely often. Our main contribution in this section is to formalize this intuition with the notion of definability (Van Den Dries, 1998; Coste, 2000; Bolte et al., 2015) and to prove Theorem 4.20, which states that for definable sa-rectangular RMDPs there always exists a stationary deterministic Blackwell optimal policy.
A concise introduction to definability. We start with the following definition. Intuitively, a set is definable if it is simple enough to be “constructed” based on polynomials, the exponential function, and canonical projections (elimination of variables).
Definition 4.11 (Definable set and definable function)
A subset of $\mathbb{R}^{n}$ is definable if it is of the form
$\{\bm{x}\in\mathbb{R}^{n}\;|\exists k\in\mathbb{N},\exists\bm{y}\in\mathbb{R}^{k},\;P\left(x_{1},...,x_{n},y_{1},...,y_{k},\exp(x_{1}),...,\exp(x_{n}),...,\exp(y_{1}),...,\exp(y_{k})\right)=0\}$ (4.3)
where $P:\mathbb{R}^{2(n+k)}\rightarrow\mathbb{R}$ is a real polynomial. For $\Omega\subset\mathbb{R}^{n}$ a definable susbet of $\mathbb{R}^{n}$, a function $f:\Omega\rightarrow\mathbb{R}^{m}$ is definable if its graph $\{(\bm{x},\bm{y})\in\Omega\times\mathbb{R}^{m}\;|\;\bm{y}=f(\bm{x})\}$ is a definable set in $\mathbb{R}^{n+m}$.
We refer to Bolte et al. (2015) and Akian et al. (2019) for concise introductions to definability and to Coste (2000) for a more in-depth treatment. It is instructive to start by studying a few simple examples.
Example 4.12 (Affine functions.)
Consider an affine function $f:\bm{x}\mapsto\bm{a}^{\top}\bm{x}+b$ for some $\bm{a}\in\mathbb{R}^{\mathcal{S}},b\in\mathbb{R}$. The graph of $f$ is $\{(\bm{x},y)\in\mathbb{R}^{\mathcal{S}}\times\mathbb{R}\;|\;\sum_{s\in\mathcal{S}}a_{s}x_{s}+b-y=0\}$, which can be written as (4.3). Therefore, affine functions are definable.
Example 4.13 ($\ell_{p}$-norms.)
Consider an $\ell_{p}$-norm for $p\in\mathbb{N}$. Then its graph is $\{(\bm{x},y)\in\mathbb{R}^{\mathcal{S}}\times\mathbb{R}\;|\;\sum_{s\in\mathcal{S}}|x_{s}|^{p}-y^{p}=0,y\geq 0\}$, which can be written as (4.3). Therefore, $\ell_{p}$-norms are definable.
Example 4.14 (Logarithm and exponential.)
It is straightforward that $x\mapsto\exp(x)$ is a definable function. The graph of the logarithm is $\{(x,y)\in\mathbb{R}_{+}^{*}\times\mathbb{R}\;|\;y=\log(x)\}$. This can be rewritten $\{(x,y)\in\mathbb{R}_{+}^{*}\times\mathbb{R}\;|\;\exp(y)=x\}$, which is a definable set. Therefore, the logarithm is a definable function.
Definable functions and definable sets are well-behaved under many useful operations, as shown in the following lemma. The proof follows directly from the definition and some properties shown in Bolte et al. (2015), and we present it for completeness in Appendix 18.
Lemma 4.15
  1. 1.
    If $A,B\subset\mathbb{R}^{n}$ are definable sets, then $A\cup B,A\cap B$ and $\mathbb{R}^{n}\setminus A$ are definable sets.
  2. 2.
    Let $f,g$ be definable functions. Then $f\circ g$, $-f$, $f+g$, $f\times g$ are definable.
  3. 3.
    For $A,B$ two definable sets and $g:A\times B\rightarrow\mathbb{R}$ a definable function, then $A\rightarrow\mathbb{R},\bm{x}\mapsto\inf_{\bm{y}\in B}g(\bm{x},\bm{y})$ and $A\rightarrow\mathbb{R},\bm{x}\mapsto\sup_{\bm{y}\in B}g(\bm{x},\bm{y})$ are definable functions.
  4. 4.
    If $A,B,g:A\rightarrow\mathbb{R}$ are definable then $g^{-1}(B)$ is definable.
Example 4.16 (Functions based on entropy.)
Let $\hat{\bm{p}}\in\Delta(\mathcal{A})$ with $\hat{p}_{s}>0,\forall\;s\in\mathcal{S}$. Consider the Kullback-Leibler divergence: $f:\bm{p}\mapsto\sum_{s\in\mathcal{S}}p_{s}\log(p_{s}/\hat{p}_{s})$ defined over $\bm{p}\in\Delta(\mathcal{S})$. Recall that $x\mapsto\log(x)$ is definable, so that $x\mapsto x\log(x)$ is also definable. By summation, the Kullback-Leibler divergence is a definable function. The case of the Burg entropy $\bm{p}\mapsto\sum_{s\in\mathcal{S}}\hat{p}_{s}\log(\hat{p}_{s}/p_{s})$ for $\bm{p}\in\Delta(\mathcal{S})$ is similar to the case of the Kullback-Leibler divergence.
The most important result for our use of definability is the following monotonicity theorem concerning definable functions of a real variable.
Theorem 4.17 (Theorem 2.1, Coste (2000))
Let $f:(a,b)\rightarrow\mathbb{R}$ be a definable function. Then there exists a finite subdivision of the interval $(a,b)$ as $a=a_{1}<a_{2}<\dots<a_{k}=b$ such that on each $(a_{i},a_{i+1})$ for $i=1,...,k-1$, $f$ is continuous and either constant or strictly monotone.
The monotonicity theorem shows that definable functions over $\mathbb{R}$ cannot oscillate infinitely often on an interval. Recall in the previous section, we have identified oscillations of the robust value functions as the main issue potentially precluding the existence of Blackwell optimal policies, see discussion after Theorem 4.4. Therefore, the monotonicity theorem will play an important role in our proof of the existence of Blackwell optimality for definable RMDPs as in Theorem 4.20.
Remark 4.18
The notion of definability is usually introduced in much more generality, e.g. (Van Den Dries, 1998; Coste, 2000; Bolte et al., 2015). We only introduce the notions necessary for this paper. For exactness, we simply note that in the literature, the notion of definability introduced in Definition 4.11 is usually referred to as definable in the real exponential field.
Definable robust MDPs. We now study sa-rectangular robust MDPs with a definable compact uncertainty set. We first note that this encompasses the vast majority of the uncertainty sets studied in the robust MDP literature. Indeed, from Lemma 4.15, we know that the set $\mathcal{U}_{sa}=\{\bm{p}\in\Delta(\mathcal{S})\;|\;d_{sa}\left(\bm{p}\right)\leq\alpha_{sa}\}$ is definable as soon as $d_{sa}:\mathbb{R}^{\mathcal{S}}\rightarrow\mathbb{R}$ is definable. From the various examples introduced above, we obtain that sa-rectangular uncertainty sets based on $\ell_{\infty}$-norm (Givan et al., 1997), $\ell_{2}$-norm (Iyengar, 2005), $\ell_{1}$-norm (Ho et al., 2021), Kullback-Leibler divergence and Burg entropy (Iyengar, 2005; Ho et al., 2022) are definable uncertainty sets.
We start by describing an appealing property of definable RMDPs: the robust value functions are themselves definable functions. In particular, we have the following proposition.
Proposition 4.19
Assume that $\mathcal{U}$ is sa-rectangular and definable. Then for any policy $\pi\in\Pi_{\sf S}$, the function $\gamma\mapsto\bm{v}^{\pi,\mathcal{U}}_{\gamma}$ is a definable function.
Proof.
Proof of Proposition 4.19. Recall that the robust value function $\bm{v}^{\pi,\mathcal{U}}_{\gamma}$ is the unique fixed-point of the following operator $\bm{v}\mapsto T^{\pi,\mathcal{U}}_{\gamma}(\bm{v})$, defined as
$T^{\pi,\mathcal{U}}_{\gamma,s}(\bm{v})=\min_{\bm{P}_{s}\in\mathcal{U}_{s}}\sum_{a\in\mathcal{A}}\pi_{sa}\bm{p}_{sa}^{\top}\left(\bm{r}_{sa}+\gamma\bm{v}\right),\forall\;s\in\mathcal{S},\forall\;\bm{v}\in\mathbb{R}^{\mathcal{S}}.$
The proof proceeds in two steps. The first step shows that $T^{\pi,\mathcal{U}}_{\gamma}$ is definable when $\mathcal{U}$ is definable. The second step shows that this is sufficient to conclude that the robust value functions are definable.
First step.
We first prove that $(\bm{v},\gamma)\mapsto T^{\pi,\mathcal{U}}_{\gamma}(\bm{v})$ is definable.
Note that a function is definable if and only if each of its components is definable (exercise 1.10, Coste (2000)). We want to show that $(\bm{v},\gamma)\mapsto\left(\sum_{a\in\mathcal{A}}\pi_{sa}\min_{\bm{p}_{sa}\in\mathcal{U}_{sa}}\bm{p}_{sa}^{\top}\left(\bm{r}_{sa}+\gamma\bm{v}\right)\right)_{s\in\mathcal{S}}$ is definable. Let $\bm{v}\in\mathbb{R}^{\mathcal{S}}$. The map $\gamma\mapsto\left(\sum_{a\in\mathcal{A}}\pi_{sa}\min_{\bm{p}_{sa}\in\mathcal{U}_{sa}}\bm{p}_{sa}^{\top}\left(\bm{r}_{sa}+\gamma\bm{v}\right)\right)_{s\in\mathcal{S}}$ is affine and therefore it is definable. We now fix $\gamma\in(0,1)$. From Lemma 4.15, if $\mathcal{U}_{sa}$ is definable for each pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, then $\bm{v}\mapsto\min_{\bm{p}_{sa}\in\mathcal{U}_{sa}}\bm{p}_{sa}^{\top}\bm{v}$ is definable, and we conclude that $\bm{v}\mapsto\sum_{a\in\mathcal{A}}\pi_{sa}\min_{\bm{p}_{sa}\in\mathcal{U}_{sa}}\bm{p}_{sa}^{\top}\left(\bm{r}_{sa}+\gamma\bm{v}\right)$ is definable for any $s\in\mathcal{S}$. This shows that $\bm{v}\mapsto(\sum_{a\in\mathcal{A}}\pi_{sa}\min_{\bm{p}_{sa}\in\mathcal{U}_{sa}}\bm{p}_{sa}^{\top}\left(\bm{r}_{sa}+\gamma\bm{v}\right))_{s\in\mathcal{S}}$ is definable.
Second step.
We now show that $\gamma\mapsto\bm{v}^{\pi,\mathcal{U}}_{\gamma}$ is definable.
[…]
Theorem 4.20
Consider a sa-rectangular robust MDP with a definable compact convex uncertainty set $\mathcal{U}$. Then there exists a stationary deterministic Blackwell optimal policy:
$\exists\;\pi\in\Pi_{\sf SD},\exists\;\gamma_{0}\in(0,1),\forall\;\gamma\in(\gamma_{0},1),\min_{\bm{P}\in\mathcal{U}}R_{\gamma}(\pi,\bm{P})\geq\sup_{\pi^{\prime}\in\Pi_{\sf S}}\min_{\bm{P}\in\mathcal{U}}R_{\gamma}(\pi^{\prime},\bm{P}).$
Proof.
[…]
The proof of Theorem 4.20 is relatively concise. It is also quite instructive, as it shows the existence of a Blackwell discount factor $\gamma_{\sf bw}\in(0,1)$, above which any discount optimal policy is also Blackwell optimal. We note that we actually proved a stronger result than the existence of a Blackwell optimal policy. In particular, we have shown the following theorem.
Theorem 4.21
Consider a sa-rectangular robust MDP with a definable compact convex uncertainty set $\mathcal{U}$. Then there exists a Blackwell discount factor $\gamma_{{\sf bw}}\in(0,1)$, such that any policy that is discount optimal for $\gamma\in(\gamma_{{\sf bw}},1)$ is also Blackwell optimal.
The existence of the Blackwell discount factor is shown in Grand-Clément and Petrik (2023) for nominal MDPs and for sa-rectangular RMDPs with polyhedral uncertainty sets. Theorem 4.21 extends this result to the larger class of definable uncertainty sets. If we know an upper bound $\hat{\gamma}<1$ on $\gamma_{{\sf bw}}$, then we can compute a Blackwell optimal policy by solving a discounted robust MDP with discount factor $\gamma=\hat{\gamma}$. We leave computing such an upper bound on $\gamma_{{\sf bw}}$ as an interesting future direction.
4.3 The case of s-rectangular robust MDPs
We now provide an instance of an s-rectangular RMDPs where there are no Blackwell optimal policies, even though the uncertainty set is definable. This may happen when all optimal policies are randomized, and the randomized optimal policies vary with the discount factor.
Proposition 4.22
There exists a s-rectangular robust MDP instance, with a compact, polyhedral uncertainty set $\mathcal{U}$, such that there is no Blackwell optimal policy:
$\forall\;\pi\in\Pi_{\sf H},\forall\;\gamma\in(0,1),\exists\;\gamma^{\prime}\in(\gamma,1),\min_{\bm{P}\in\mathcal{U}}R_{\gamma^{\prime}}(\pi,\bm{P})<\sup_{\pi^{\prime}\in\Pi_{\sf S}}\min_{\bm{P}\in\mathcal{U}}R_{\gamma^{\prime}}(\pi^{\prime},\bm{P}).$
Proof.
Proof.
In Figure 3, we adapt the example from Figure 3 in Wiesemann et al. (2013). There are three states and two actions.
(a) Transition for action $a_{1}$
(b) Transition for action $a_{2}$
Figure 3: Transitions and rewards for a s-rectangular RMDP instance with no Blackwell optimal policy.
[…]
We note that the counterexample in the proof of Proposition 4.22 is much simpler than the counterexample from Section 4.2.1 for sa-rectangular RMDPs. In particular, Proposition 4.22 shows that Blackwell optimal policies may fail to exist for s-rectangular RMDPs, even in the simple setting where the uncertainty set is a polytope parametrized by $p\in[0,1]$. As a possible next step, it sounds promising to study the existence and tractability of $\epsilon$-Blackwell optimal policies for s-rectangular RMDPs.
5 Algorithms
In this section, we discuss various iterative algorithms to compute average optimal and Blackwell optimal policies. Our results for s-rectangular RMDPs from Section 3.3 and Section 4.3 suggest that it may be difficult to compute average and Blackwell optimal policies in this case. Therefore, we focus on sa-rectangular robust MDPs. Since for sa-rectangular RMDPs an average optimal policy may be chosen stationary and deterministic, we define the optimal gain $\bm{g}^{\star}\in\mathbb{R}^{\mathcal{S}}$ as
$g^{\star}_{s}=\max_{\pi\in\Pi_{\sf SD}}\inf_{\bm{P}\in\mathcal{U}}\lim_{T\rightarrow+\infty}\mathbb{E}_{\pi,\bm{P}}\left[\frac{1}{T+1}\sum_{t=0}^{T}r_{s_{t}a_{t}s_{t+1}}\;|\;s_{0}=s\right],\forall\;s\in\mathcal{S}.$ (5.1)
We now introduce three algorithms to compute the optimal gain $\bm{g}^{\star}$ when $\mathcal{U}$ is sa-rectangular and definable.
5.1 Algorithms based on the Blackwell discount factor
As discussed in Section 4.2.3, when $\mathcal{U}$ is definable, we have shown that there exists a discount factor $\gamma_{{\sf bw}}\in(0,1)$, such that any policy that is $\gamma$-discount optimal for $\gamma\in(\gamma_{{\sf bw}},1)$ is also Blackwell optimal. Therefore, we can compute Blackwell optimal and average optimal policies by solving a sequence of discounted RMDPs with discount factors increasing to $1$. An immediate consequence is the following theorem.
Theorem 5.1
Let $\mathcal{U}$ be a definable compact convex uncertainty set and let $\left(\bm{v}^{t}\right)_{t\geq 1}$ and $\left(\pi^{t}\right)_{t\in\mathbb{N}}$ be the iterates of Algorithm 1. Then $\left(\bm{v}^{t}\right)_{t\geq 1}$ converges to $\bm{g}^{\star}\in\mathbb{R}^{\mathcal{S}}$, and $\pi^{t}$ is both Blackwell optimal and average optimal for $t\in\mathbb{N}$ large enough.
Unfortunately, computing the Blackwell discount factor $\gamma_{{\sf bw}}$ appears challenging, so that we are not able to provide a convergence rate for Algorithm 1. When $\mathcal{U}$ is polyhedral, Grand-Clément and Petrik (2023) obtains an upper bound on $\gamma_{{\sf bw}}$, but this upper bound is too close to $1$ to be of practical use. We also note that Algorithm 1 requires solving a robust MDP at every iteration, which may be computationally intensive. We now describe two algorithms with a lower per-iteration complexity.
Algorithm 1 Iterative algorithm for average optimality: limit of discounted values
Input: An increasing sequence of discount factors $\left(\gamma_{t}\right)_{t\in\mathbb{N}}$ with $\lim_{t\rightarrow+\infty}\gamma_{t}=1$.
Initialize $\bm{v}^{0}=\bm{0}\in\mathbb{R}^{\mathcal{S}},\pi^{0}\in\Pi_{\sf S}$.
for $t\in\mathbb{N}$ do
     $\bm{v}^{t+1}=(1-\gamma_{t})\bm{v}^{\star,\mathcal{U}}_{\gamma_{t}}$
     $\pi^{t+1}\in\Pi_{\sf S}$ such that $(1-\gamma_{t})\bm{v}^{\pi^{t+1},\mathcal{U}}_{\gamma_{t}}=\bm{v}^{t+1}$
end for
5.2 Algorithms based on value iteration
We now introduce two value iteration algorithms to compute the optimal worst-case average return when the uncertainty set $\mathcal{U}$ is definable.
Algorithm based on increasing horizon.
Algorithm 2 Iterative algorithm for average optimality: increasing horizon
Initialize $\bm{v}^{0}=\bm{0}\in\mathbb{R}^{\mathcal{S}}$.
for $t\in\mathbb{N}$ do
     $v^{t+1}_{s}=\max_{a\in\mathcal{A}}\min_{\bm{p}_{sa}\in\mathcal{U}_{sa}}\bm{p}_{sa}^{\top}\left(\bm{r}_{sa}+\bm{v}^{t}\right),\forall\;s\in\mathcal{S}$.
end for
We start with Algorithm 2, which computes the optimal value functions of a finite-horizon RMDP as the horizon increases to +$\infty$+ ∞. The following theorem shows that Algorithm 2 is correct.
Theorem 5.2
Let $\mathcal{U}$ be a definable compact convex uncertainty set and let $\left(\bm{v}^{t}\right)_{t\geq 1}$ be the iterates of Algorithm 2. Then $\left(\frac{1}{t}\bm{v}^{t}\right)_{t\geq 1}$ converges to $\bm{g}^{\star}\in\mathbb{R}^{\mathcal{S}}$.
Proof.
[…]
The proof of Theorem 5.2 is relatively short. This conciseness comes from invoking powerful results from the stochastic game literature (Bolte et al., 2015).
Algorithm based on increasing discount factor.
Algorithm 3 Iterative algorithm for average optimality: increasing discount factor
Input: an increasing sequence $\left(\omega_{t}\right)_{t\geq 0}\in\mathbb{R}^{\mathbb{N}}_{+}$ with $\lim_{t\rightarrow+\infty}\omega_{t}=+\infty,\lim_{t\rightarrow+\infty}\omega_{t}/\omega_{t+1}=1$.
Initialize $\bm{v}^{0}=\bm{0}\in\mathbb{R}^{\mathcal{S}}$.
for $t\in\mathbb{N}$ do
     $\gamma_{t}\;=\;\frac{\omega_{t}}{\omega_{t+1}}$
     $v^{t+1}_{s}\;=\;\max_{a\in\mathcal{A}}\min_{\bm{p}_{sa}\in\mathcal{U}_{sa}}\bm{p}_{sa}^{\top}\left((1-\gamma_{t})\bm{r}_{sa}+\gamma_{t}\bm{v}^{t}\right),\forall\;s\in\mathcal{S}$.
end for
We now consider Algorithm 3, which builds upon value iteration for RMDPs with an increasing sequence of discount factors. We have the following theorem.
Theorem 5.3
Let $\mathcal{U}$ be a definable compact convex uncertainty set and let $\left(\bm{v}^{t}\right)_{t\geq 1}$ be the iterates of Algorithm 3. Then $\left(\bm{v}^{t}\right)_{t\geq 1}$ converges to $\bm{g}^{\star}\in\mathbb{R}^{\mathcal{S}}$.
The proof of Theorem 5.3 is deferred to Appendix 20. We only provide the main steps here. The first step is to show the following lemma, which decomposes the suboptimality gap of the current iterates of Algorithm 1.
Lemma 5.4
There exist a period $t_{0}\in\mathbb{N}$ and two sequences of non-negative scalars $\left(\epsilon_{t}\right)_{t\geq t_{0}}$ and $\left(\delta_{t}\right)_{t\geq t_{0}}$ such that
$\|\bm{v}^{t}-\bm{g}^{\star}\|_{\infty}\leq\epsilon_{t}+\delta_{t},\forall\;t\geq t_{0},$
with $\lim_{t\rightarrow 0}\epsilon_{t}\rightarrow 0$ and with, for $t\geq t_{0}$ and $\pi^{B}$ a Blackwell optimal policy,
$\delta_{t+1}$ $=\gamma_{t}\delta_{t}+e_{t},$ (5.2)
$e_{t}$ $=\|(1-\gamma_{t+1})\bm{v}^{\pi^{B},\mathcal{U}}_{\gamma_{t+1}}-(1-\gamma_{t})\bm{v}^{\pi^{B},\mathcal{U}}_{\gamma_{t}}\|_{\infty}.$ (5.3)
The difficult part of the proof is to prove that the sequence $\left(\delta_{t}\right)_{t\geq t_{0}}$ converges to 00. To show $\lim_{t\rightarrow+\infty}\delta_{t}\rightarrow 0$, we use the inductive step (5.2) and $\gamma_{t}=\omega_{t}/\omega_{t+1}$ to show
$\delta_{t}=\frac{1}{\omega_{t+1}}\sum_{j=t_{0}}^{t}\omega_{j+1}e_{j},\forall\;t\geq t_{0}.$ (5.4)
Equation (5.4) reveals that the expression of $\delta_{t}$ resembles a weighted Cesaro average of the sequence $\left(e_{t}\right)_{t\geq t_{0}}$. It is known that the Cesaro average of a converging sequence also converges to the same limit; we prove Theorem 5.3 by extending this result to weighted Cesaro averages and by invoking existing results from SGs to obtain that $\lim_{t\rightarrow+\infty}e_{t}=0.$
Comparison with previous work. We now compare Theorem 5.3 with related results from the RMDP literature. The authors in Tewari and Bartlett (2007) prove Theorem 5.3 with $\omega_{t}=t+1,$ i.e., with $\gamma_{t}=(t+1)/(t+2)$ and for the case of sa-rectangular uncertainty sets with $\ell_{\infty}$-balls. Tewari and Bartlett (2007) show the Lipschitz continuity of the robust value functions to simplify the term $e_{t}$ from Lemma 5.4. In particular, for sa-rectangular polyhedral $\mathcal{U}$, for $t$ large enough, there exists a constant $L>0$ such that
$e_{t}=\|(1-\gamma_{t+1})\bm{v}^{\pi^{B},\mathcal{U}}_{\gamma_{t+1}}-(1-\gamma_{t})\bm{v}^{\pi^{B},\mathcal{U}}_{\gamma_{t}}\|_{\infty}\leq L\cdot\left(\gamma_{t+1}-\gamma_{t}\right).$ (5.5)
Once (5.5) is established, one can conclude that
$\delta_{t+1}=\gamma_{t}\delta_{t}+L\cdot\left(\gamma_{t+1}-\gamma_{t}\right)$ (5.6)
and Lemma 8 in Tewari and Bartlett (2007) shows that a sequence $\left(\delta_{t}\right)_{t\geq 0}$ satisfying (5.6) converges to 00.
The authors in Wang et al. (2023) consider Algorithm 3 with $\gamma_{t}=(t+1)/(t+2)$ for the case of general sa-rectangular uncertainty set (potentially non-polyhedral), with the assumption that the Markov chain associated with any transition probabilities and any policy is unichain. The analysis of Algorithm 3 in Wang et al. (2023) follows the same lines as the proof in Tewari and Bartlett (2007), and the authors conclude by proving that the Lipschitzness property (5.5) holds for unichain sa-rectangular RMDPs. However, we do not assume that the RMDP is unichain, and in all generality the normalized robust value function $\gamma\mapsto(1-\gamma)\bm{v}^{\pi,\mathcal{U}}_{\gamma}$ may not be Lipschitz continuous for a fixed $\pi\in\Pi_{\sf S}$, as we prove in Appendix 19. We sidestep this issue by leveraging existing results from SGs (Bolte et al., 2015) and elementary results from real analysis.
Stopping criterion and complexity. We have proved that Algorithm 1, Algorithm 2, and Algorithm 3 converge to an optimal gain. However, we would like to emphasize that we do not introduce a stopping criterion for these algorithms. Even for the simpler case of nominal MDPs, in all generality, there is no known stopping criterion for value iteration for the average return criterion, e.g. Ashok et al. (2017) and section 9.4 in Puterman (2014). We leave finding a stopping criterion as an open question in this paper, noting that this is not addressed in previous papers on average-reward RMDPs (Tewari and Bartlett, 2007; Wang et al., 2023).
We also comment on the complexity of computing an average-optimal policy (and not just computing an optimal gain). For sa-rectangular RMDPs, this may be a difficult task. Indeed, we described the connections between sa-rectangular RMDPs with average return and mean-payoff perfect information stochastic games in Section 2.2. There is no known polynomial-time algorithm for solving this class of SGs (Condon, 1992; Zwick and Paterson, 1996; Andersson and Miltersen, 2009) even after more than six decades since their introduction in the game theory literature (Gillette, 1957), and this is regarded as one of the major open questions in algorithmic game theory. For this reason, in this paper, we have focused on the properties of average and Blackwell optimal policies and we leave designing polynomial-time algorithms for solving average return sa-rectangular RMDPs as an open question.
6 Numerical experiments
In this section, we compare the empirical performances of the algorithms introduced in Section 5. Our goal is to test the practical convergence of Algorithm 1, Algorithm 2 and Algorithm 3 on various sa-rectangular RMDPs instances.
Test instances. We consider three different nominal MDP instances: a machine replacement problem (Delage and Mannor, 2010; Wiesemann et al., 2013), a forest management instance (Possingham and Tuck, 1997; Cordwell et al., 2015), and an instance inspired from healthcare (Goyal and Grand-Clément, 2022). In the machine replacement problem, the goal is to compute a replacement and repair schedule for a line of machines. In the forest management instance, a forest grows at every period, and the goal is to balance the revenue from wood cutting and the risk of wildfire. In the healthcare instance, the goal is to plan the treatment of a patient, avoiding the mortality state while reducing the invasiveness of the treatment. We refer to the nominal transition probabilities in these instances as $\bm{P}_{\sf nom}$ and we represent the RMDP instances in Appendix 21.1.
Construction of the uncertainty sets. For each of these three RMDP instances, we consider two types of uncertainty sets. The first uncertainty set is based on box inequalities:
$\mathcal{U}_{sa}^{\sf box}=\{\bm{p}\in\Delta(\mathcal{S})\;|\;\bm{p}_{sa}^{\sf low}\leq\bm{p}\leq\bm{p}_{sa}^{\sf up}\},\forall\;(s,a)\in\mathcal{S}\times\mathcal{A}$
with $\bm{p}_{sa}^{\sf low},\bm{p}_{sa}^{\sf up}\in[0,1]^{\mathcal{S}}$ two vectors such that $\bm{p}_{sa}^{\sf low}\leq\bm{p}_{{\sf nom},sa}\leq\bm{p}_{sa}^{\sf up}$ for $(s,a)\in\mathcal{S}\times\mathcal{A}$. The second uncertainty set is based on ellipsoid constraints:
$\mathcal{U}_{sa}^{\sf\ell_{2}}=\{\bm{p}\in\Delta(\mathcal{S})\;|\;\|\bm{p}-\bm{p}_{{\sf nom},sa}\|_{2}\leq\alpha\},\forall\;(s,a)\in\mathcal{S}\times\mathcal{A}$
where $\alpha>0$ is a scalar. Box inequalities have been used in applications of RMDPs in healthcare (Goh et al., 2018), while uncertainty based on $\ell_{2}$-distance can be seen as conservative approximations of sets based on relative entropy (Iyengar, 2005). Note that both $\mathcal{U}^{\ell_{2}}$ and $\mathcal{U}^{\sf box}$ are definable sets, with $\mathcal{U}^{\sf box}$ being polyhedral. Additionally, there exist efficient algorithms to evaluate the robust Bellman operator as in (2.3) for $\mathcal{U}^{\sf box}$ (e.g. proposition 3 in Goh et al. (2018)). For $\mathcal{U}^{\sf\ell_{2}}$, evaluating the robust Bellman operator requires solving a convex program. Still, we can obtain a closed-form expression assuming that $\alpha>0$ is sufficiently small, as we detail in Appendix 21.2. We choose a uniform initial distribution for all instances.
Empirical setup. We run Algorithm 1, Algorithm 2 and Algorithm 3 for $T=50$ iterations, on the Machine, Forest, and Healthcare RMDP instances with $S=100$ states. For $\mathcal{U}^{\sf\ell_{2}}$ we choose a radius of $\alpha=0.05$ and for $\mathcal{U}^{\sf box}$ we choose the upper and lower bound on each coefficient to allow for $5\%$ deviations from the nominal distributions. For Algorithm 3, we choose $\omega_{t}=t+1$. We provide more details in Appendix 21.2. To implement Algorithm 1, we need to solve a robust MDP at every iteration. To do so, we implement the two-player strategy iteration (Hansen et al., 2013) with warm-starts, see more details in Appendix 21.3. We implement all algorithms in Python 3.8.8.
Numerical results. We first present the convergence rate as a function of the number of iterations of all our algorithms in Figure 4 for $\mathcal{U}^{\sf\ell_{2}}$ and in Figure 5 for $\mathcal{U}^{\sf box}$. Algorithm 1, Algorithm 2 and Algorithm 3 appear to have comparable convergence speed on our three RMDP instances, despite Algorithm 1 solving a discounted RMDP at each iteration, compared to Algorithm 2 and Algorithm 3 which only require to evaluate the robust Bellman operator. We also present the total computation time as a function of the number of iterations in Figure 6-7. We note that Algorithm 1 is much slower than Algorithm 2 and Algorithm 3, even with warm-start. Indeed, at every iteration, Algorithm 1 requires running two-player strategy iteration for solving a discounted RMDP, whereas the other two algorithms only require evaluating the Bellman operator. We also note that running our algorithms for $\mathcal{U}^{\sf\ell_{2}}$ is faster than for $\mathcal{U}^{\sf box}$, since the Bellman update for $\mathcal{U}^{\sf box}$ requires to sort a vector of size $|\mathcal{S}|$, and then compute a maximum, whereas for $\mathcal{U}^{\sf\ell_{2}}$ we only need to compute a maximum, see the implementation details in Appendix 21.2.
(a) Machine
(b) Healthcare
(c) Forest
Figure 4: Convergence to the optimal average return for various RMDP instances, for $\mathcal{U}^{\sf\ell_{2}}$ based on ellipsoids.
(a) Machine
(b) Healthcare
(c) Forest
Figure 5: Convergence to the optimal average return for various RMDP instances, for $\mathcal{U}^{\sf box}$ based on box inequalities.
(a) Machine
(b) Healthcare
(c) Forest
Figure 6: Total computation time as a function of the number of iterations for various RMDP instances, for $\mathcal{U}^{\sf\ell_{2}}$ based on ellipsoids.
(a) Machine
(b) Healthcare
(c) Forest
Figure 7: Total computation time as a function of the number of iterations for various RMDP instances, for $\mathcal{U}^{\sf box}$ based on box inequalities.
7 Conclusion
Our paper addresses several important issues in the existing literature and derives the fundamental properties of average optimal and Blackwell optimal policies. In particular, our work highlights important distinctions between the widely-studied framework of discounted RMDPs and the less-studied frameworks of RMDPs with average optimality and Blackwell optimality. We view the non-existence of stationary average optimal policies for s-rectangular RMDPs and the non-existence of Blackwell optimal policies for sa-rectangular RMDPs (in all generality) as surprising results. The notion of definable uncertainty sets is also of independent interest, characterizing the cases where the value functions are well-behaved and iterative algorithms asymptotically converge to the optimal gain in the sa-rectangular case. Finally, our work opens new research avenues for RMDPs. Among them, deriving an efficient algorithm for computing an average optimal policy for sa-rectangular RMDPs appears crucial, but this may be difficult since a similar question remains open after several decades in the literature on SGs. Important other research questions include studying in more detail the particular case of irreducible instances, which considerably simplifies the main technical challenges, and studying the case of distributionally robust MDPs. Understanding if Markovian policies are sufficient for average return s-rectangular RMDPs also looks like a promising next direction.